Find the area enclosed by the curve r = 5cosθ


1
Expert's answer
2021-12-27T04:03:17-0500

Solution:


Assuming that we are dealing with polar coordinates, the curve :


"r=5cos\theta"


for 0 < θ < π traces out a circle of radius "\frac{5}{2}" centered at "x=\frac{5}{2}" , y=0. For π < θ < 2π the same circle is repeated. The area within the circle is, therefore,



"\pi(\frac{5}{2})^2=\frac{25\pi}{4}"